Linda estimates that she spends 1/3 of her income on housing and utilities and 1/4 of her income on food. What fraction of her income is left after she covers these expenses?
Answer by nerdybill(7384)

Linda estimates that she spends 1/3 of her income on housing and utilities and 1/4 of her income on food. What fraction of her income is left after she covers these expenses?
1/3 + 1/4 + x = 1
multiplying both sides by 12 to get rid of denominators:
4 + 3 + 12x = 12
7 + 12x = 12
12x = 5
x = 5/12